Perdue Company purchased equipment on April 1 for $43,470. The equipment was expected to have a useful life of three years, or 7,020 operating hours, and a residual value of $1,350. The equipment was used for 1,300 hours during Year 1, 2,500 hours in Year 2, 2,100 hours in Year 3, and 1,120 hours in Year 4. Determine the amount of depreciation expense for the years ended December 31, Year 1, Year 2, Year 3, and Year 4, by (a) the straight-line method, (b) units-of-activity method, and (c) the double-declining-balance method. Note: FOR DECLINING BALANCE ONLY, round the multiplier to four decimal places. Then round the answer for each year to the nearest whole dollar.
Answer #1

(A)

Under straight line method of depreciation,

Depreciation per annum = (initial cost - salvage value)/useful life

= ($43470 - $1350)/3

= $14040

For year 1:

Depreciation expenses = $14040 x (9/12) = $10530

For year 2:

Depreciation expenses = $14040

For year 3:

Depreciation expenses = $14040

for year 4:

depreciation expense = $14040 x (3/12) = $3510

(B)

Under units of activity method,

depreciable value = initial cost - salvage value

= $43470 - $1350 = $42120

Depreciation per annum = depreciable value x (units produced during the year/estimated total activity units)

Therefore,

For year 1,

Depreciation = $42120 x (1300/7020) = $7800

For year 2,

Depreciation = $42120 x (2500/7020) = $15000

For year 3,

Depreciation = $42120 x (2100/7020) = $12600

For year 4,

Depreciation = $42120 x (1120/7020) = $6720